OpenStudy (anonymous):

A line has a slope of -1.5 and a y-intercept of (0,3). What is the value of y when x=6?

OpenStudy (lalaly):

the equation is y-3=-1.5x y=-1.5x+3 substitute x=6 y=-1.5(6)+3 = -9+3= -6